OpenStudy (anonymous):

Solving with a matrix, once again. Enter the answer as an ordered pair. 4x - 4y = 12 -7x + 4y = -36 All of this confuses me..

OpenStudy (anonymous):

if you add the second line to the first, what do you get ?

OpenStudy (anonymous):

Should I cancel both of the fours?

OpenStudy (anonymous):

yes, that's the point. if you add those two lines you get 4x -7x -4y +4y = 12 - 36 and -4y + 4y = 0 so you have 4x-7x = 12-36

OpenStudy (anonymous):

\[A = \left[\begin{matrix}4 & -4 \\ -7 & 4\end{matrix}\right], B = \left[\begin{matrix}12 \\ -36 \end{matrix}\right]\]\[A ^{-1} = 1/12\left[\begin{matrix}-4 & -4 \\ -7 & -4\end{matrix}\right]\]\[A ^{-1} * \left[\begin{matrix}12 \\ -36\end{matrix}\right] = [-16,-19]\] x = -16 y = -19 Check: 4(-16) - 4(-19) = 12 -7(-16) + 4(19) = -36

OpenStudy (anonymous):

To calculate the inverse matrix of a 2 by 2 matrix: \[K = \left[\begin{matrix}a & b \\ c & d\end{matrix}\right], K ^{-1} = \left[\begin{matrix}d/(ad-bc) & -b/(ad-bc) \\ -c/(ad-bc) & a/(ad-bc)\end{matrix}\right]\]

OpenStudy (anonymous):

To obtain the answer you then just multiply the constant matrix B by the inverse matrice of A